If the exterior angles of a convex octagon are x+13, 2x-1, 2x+6, 2x+12, 2x-17, 3x-4, 3x-10, and 4x, calculate the smallest of the eight angles.
choices
A) 19 degrees
B) 21 degrees
C) 32 degrees
D) 37 degrees

Answer by stanbon(75887)

If the exterior angles of a convex octagon are x+13, 2x-1, 2x+6, 2x+12, 2x-17, 3x-4, 3x-10, and 4x, calculate the smallest of the eight angles.
The sum of these angles must be 360
EQUATION
x+13+2x-1+2x+6+2x+12+2x-17+3x-4+3x-10+4x=360
19x-1=360
19x=361
x=19
2x-17=21 degrees (smallest of the eight angles)
Cheers,
Stan H.